# Improving Hospital Risk Prediction with Knowledge-Augmented Multimodal EHR Modeling Improving Hospital Risk Prediction with Knowledge-Augmented Multimodal EHR Modeling

### v1

@@ 3 Background
Problem statements. To demonstrate the effectiveness of our method, we study two popular clinical tasks: in-hospital mortality prediction and 30-day readmission prediction, e.g., (Jiang et al. 2025; Cai et al. 2016). We define these problems formally after introducing some notation.
@@
30-day Readmission Problem: The objective is to determine whether the patient is readmitted to the hospital within 30 days following discharge from visit . We define the readmission indicator as:
@@
4 KAMELEON Framework
We propose a hybrid framework, KAMELEON, that integrates multimodal EHR data, including structured clinical components and unstructured physician notes (we will provide more details in Section 5.1), and external biomedical knowledge to predict two key clinical outcomes: in-hospital mortality and 30-day readmission. As shown in Figure 1, KAMELEON consists of two components: (1) an unstructured encoder (Section 4.1) that processes clinical notes and retrieves relevant biomedical knowledge using a PubMed-derived graph and knowledge-augmented reasoning, and outputs a prediction, along with its reasoning; and (2) a structured encoder (Section 4.2) that combines multiple time-series corresponding to vitals and tabular datasets (labs, medications, etc), along with the outputs from (i.e., both the prediction and the embedding associated with the reasoning it produces) with static features for downstream prediction. The notations used in Section 4, Algorithm 1, and 2 are explained in Table 1.
4.1 Unstructured Text Encoder,
@@
Generating Context: The first step of the framework is generating patient context. We use EHR data, including physicians’ notes, patient conditions, prescribed medications, and procedures. Since physician notes can be lengthy and exceed the context window of a small, locally fine-tuned LLM, we summarize them using an LLM. This approach addresses Challenge 2 (Long Context).
@@
Knowledge Graph Retrieval: To introduce domain-level medical knowledge, we build a biomedical knowledge graph (KG) by combining the Unified Medical Language System (UMLS)(Bodenreider 2004), PubMed abstracts, and large language model (LLM)-generated entity-relation triples. UMLS provides standardized biomedical concepts and relationships, and the entity-relation triples are structured facts extracted by the LLM in the form , capturing semantic connections. We apply the Leiden algorithm for community detection (Traag, Waltman, and Eck 2019), which partitions the KG into semantically coherent and well-connected subgraphs. After clustering, we employ a separate LLM to generate a textual summary for each cluster. These summaries are produced by reasoning over the relationships among entities within each cluster, capturing the latent biomedical semantics encoded in the graph. Each summary serves as a high-level abstraction of the biomedical concepts and interactions within a subgraph. We embed these summaries using SentenceTransformer (MiniLM-L6-v2) (Reimers and Gurevych 2019) and retrieve the most relevant ones for each patient by computing semantic similarity with the embedded patient context. This process directly addresses Challenge 3 (Specialized Medical Domains) by enriching patient context with structured, domain-specific knowledge, improving the model’s understanding of specialized medical terminology.
@@
Finding Similar Patients: We provide additional context, by retrieving semantically similar patient visits using precomputed visit-level embeddings and a similarity index using the FAISS (Douze et al. 2024) library with inner-product search on L2-normalized embeddings, which effectively approximates cosine similarity. For each target patient visit, we retrieve the top 50 most similar patients while excluding self-matches and other visits from the same individual. Each retrieved patient is scored by similarity, and we filter them into positive and negative cohorts based on matching or non-matching ground truth labels (e.g., readmission vs. no readmission). The final output includes the top-k positive and negative similar patients (with k=1,2). Unlike KARE (Jiang et al. 2025), we also provide the physician notes of the retrieved similar patients, enabling the language model to leverage more clinical context when assessing patient risk.
@@
Reasoning Module: In this module, we prepare inputs to fine-tune the LLM for clinical prediction. For each patient visit, we create a prompt with the patient’s context with the top-k most similar cases retrieved earlier. These similar cases guide the model by highlighting patterns in clinically comparable scenarios. We also add biomedical knowledge summaries from clustered subgraphs of a PubMed knowledge graph, providing literature-based context. Combining patient data, historical cases, and domain knowledge, we fine-tune the LLM to produce task-specific predictions with interpretable reasoning, supporting each outcome.
@@
Finetuning LLM: We fine-tune a LLaMA-3 8B model using the Unsloth framework (Touvron et al. 2024; Daniel Han and team 2023), which enables memory-efficient training via 4-bit quantization and Low-Rank Adaptation (LoRA) (Hu et al. 2022). Prediction tasks are framed as instruction-following using Alpaca-style prompts with task description, patient context, and optional justification. Each prompt combines clinical notes, retrieved similar cases, and biomedical knowledge summaries. The model is trained via supervised learning to generate both predictions and reasoning. We use limited training steps with gradient accumulation and sequence lengths up to tokens. Unlike KARE, which trains larger models with higher compute, our method uses smaller, quantized models to reduce computational cost while maintaining interpretability and performance. Algorithm 1 outlines the training and inference procedures of . Additionally, an example prompt and its overall structure are provided in the Appendix Section A.3.
4.2 Structured Data Encoder,
@@
Time-varying variables: We extract hourly vitals and diagnoses in MIMIC-III, including heart rate, systolic and diastolic blood pressure, mean blood pressure, oxygen saturation, Glasgow Coma Scale (GCS) scores, glucose level, respiratory rate, temperature, weight, and pH.
Static metadata: Demographic and admission features include gender, ethnicity, admission type, location, insurance, language, religion, and hospital outcome.
Diagnoses, procedures, medications: ICD-9 codes, drug names are encoded via one-hot or counts. We compute binary indicators for key comorbidities (e.g., sepsis, infection, organ failure, dementia, cancer, diabetes).
Structured Data Preprocessing: The structured data is first transformed using a discretization step to enforce uniform temporal resolution and impute missing values. This is followed by normalization using precomputed mean and standard deviation statistics over the continuous variables.
Incorporating LLM Output: To augment the structured input, we include the LLM’s prediction and its tokenized reasoning. A Word2Vec (Mikolov et al. 2013) model is trained on this reasoning to generate embeddings. Each paragraph is represented by averaging its word vectors, producing a fixed-size vector. This reasoning vector is concatenated with the output label to form a feature vector for each visit.
@@ 4.3 Training
In our experiments, we benchmark several ML models for as KAMELEON-X, where X represents logistic regression, random forests, LSTM, LightGBM, MLP, or XGBoost, selected for its effectiveness in capturing clinical patterns. For MLP, we use weighted binary cross-entropy loss
@@
where and are the positive and negative class weights, respectively, used to address class imbalance by giving more emphasis to the minority class. We further use Synthetic Minority Over-sampling Technique (SMOTE) (Chawla et al. 2002) to mitigate class imbalance, addressing Challenge 4 (Highly imbalanced data). The pipeline for is in Algorithm 2.
| Model | Acc | Prec0 | Prec1 | Rec0 | Rec1 | Macro F1 | AUROC | AUPRC | ||
| Task: 30-Day Readmission Prediction | ||||||||||
| LogReg (C=1.0) (Harutyunyan et al. 2019) | ✓ | – | 0.831 | 0.869 | 0.036 | 0.951 | 0.013 | 0.464 | 0.463 | 0.090 |
| MLP (Bopche et al. 2024) | ✓ | – | 0.828 | 0.876 | 0.182 | 0.934 | 0.100 | 0.516 | 0.559 | 0.165 |
| LSTM (Harutyunyan et al. 2019) | ✓ | – | 0.820 | 0.876 | 0.163 | 0.925 | 0.100 | 0.512 | 0.569 | 0.152 |
| Claude-3.7-Sonnet | – | ✓ | 0.240 | 0.790 | 0.199 | 0.068 | 0.927 | 0.227 | 0.498 | 0.199 |
| Medgemma-4b-it | – | ✓ | 0.350 | 0.770 | 0.190 | 0.270 | 0.690 | 0.350 | 0.480 | 0.190 |
| Llama3-Med42-8B | – | ✓ | 0.390 | 0.800 | 0.210 | 0.360 | 0.670 | 0.410 | 0.510 | 0.210 |
| KARE | – | ✓ | 0.660 | 0.870 | 0.130 | 0.720 | 0.280 | 0.480 | 0.506 | 0.195 |
| Task: In-Hospital Mortality Prediction | ||||||||||
| Logistic Regression (Harutyunyan et al. 2019) | ✓ | – | 0.850 | 0.912 | 0.340 | 0.916 | 0.331 | 0.625 | 0.624 | 0.190 |
| LSTM (Harutyunyan et al. 2019) | ✓ | – | 0.690 | 0.800 | 0.260 | 0.800 | 0.250 | 0.530 | 0.560 | 0.240 |
| BalancedRF | ✓ | – | 0.810 | 0.950 | 0.340 | 0.820 | 0.700 | 0.670 | 0.860 | 0.475 |
| LightGBM (Bopche et al. 2024) | ✓ | – | 0.890 | 0.930 | 0.510 | 0.940 | 0.480 | 0.720 | 0.866 | 0.534 |
| MLP (Bopche et al. 2024) | ✓ | – | 0.870 | 0.920 | 0.430 | 0.920 | 0.430 | 0.680 | 0.829 | 0.426 |
| Claude-3.7-Sonnet | – | ✓ | 0.800 | 0.890 | 0.120 | 0.880 | 0.130 | 0.510 | 0.510 | 0.110 |
| Medgemma-4b-it | – | ✓ | 0.120 | 0.950 | 0.100 | 0.020 | 0.990 | 0.120 | 0.510 | 0.110 |
| Llama3-Med42-8B | – | ✓ | 0.160 | 0.950 | 0.120 | 0.100 | 0.970 | 0.190 | 0.530 | 0.120 |
| KARE | – | ✓ | 0.614 | 0.890 | 0.134 | 0.641 | 0.413 | 0.474 | 0.527 | 0.125 |
| KAMELEON-BalancedRF | ✓ | ✓ | 0.880 | 0.930 | 0.490 | 0.934 | 0.492 | 0.710 | 0.876 | 0.543 |
5 Results
We evaluate our model on the two clinical prediction tasks: in-hospital mortality and 30-day readmission. Our experiments compare performance against strong baselines, including general and medical LLMs, and assess the value of key model components through ablation studies. Some of our results are discussed in Section B Appendix.
5.1 Experimental Setup
Datasets. We use the MIMIC-III dataset (Johnson et al. 2016), which includes structured and unstructured data for over 40,000 ICU patients. It includes structured data (demographics, vitals, labs, admissions, ICD-9-CM codes) and unstructured clinical text (physician notes, discharge summaries, radiology reports). For this study, we focus on physician-authored notes containing clinical reasoning, assessments, and treatment plans. Here, we exclude discharge summaries and notes written after outcomes to prevent label leakage. Only of notes mention hospice, indicating rare explicit terminal indicators.
Biomedical Knowledge. We use abstracts from the annual PubMed Baseline dataset, comprising over 36 million biomedical citation records, to build a medical knowledge graph that enriches LLM input, and reduces hallucinations. We also incorporate UMLS (Bodenreider 2004) to construct concept-centric subgraphs from EHR data.
Baselines. We compare against Claude 3.5 Sonnet (Anthropic 2024), MedGemma (Sellergren et al. 2025), LLaMA3-Med (Christophe et al. 2024), and KARE (Jiang et al. 2025), as well as structured-data models including logistic regression, tree-based models, and MLPs used in prior work on MIMIC III (Harutyunyan et al. 2019; Lundberg et al. 2020; Purushotham et al. 2018). All LLMs are evaluated in a zero-shot setting with the same patient-context prompt. KARE uses similar patient retrieval but lacks clinical notes. Our model incorporates retrieved notes for better context. Implementation details are in the Appendix.
Metrics. We report AUROC, AUPRC, accuracy, precision, recall, and macro F1-score. Metrics are computed on a held-out test set with similar standardized preprocessing.
5.2 Performance of KAMELEON
30-Day Readmission Prediction:
As shown in Table 2, our framework with Balanced Random Forest (KAMELEON-BalancedRF) classifier achieves the highest AUROC and notably improves recall on positive cases to , a crucial metric since identifying patients at risk of readmission is clinically imperative. The KAMELEON-MLP model, while achieving the highest overall accuracy and macro F1 , still attains a recall of on positive cases, illustrating the persistent challenge in detecting rare events.
Unstructured LLM-based baselines such as Claude-3.7-Sonnet, MedGemma, LLaMA3-Med, and KARE show substantially lower recall for positives (below 0.3), suggesting that these models struggle to identify the minority class without further fine-tuning or domain-specific adaptation.
While prior studies like Morgan et al. (2019) (Morgan et al. 2019) reported AUROCs up to 0.81 for readmission, and general models typically ranged from 0.61 to 0.73 (Dhalluin et al. 2020; Matheny et al. 2021), our multi-modal approach effectively captures complex clinical nuances.
In-Hospital Mortality Prediction: Mortality prediction is less imbalanced, with approximately positive cases. Table 2 reports that KAMELEON-XGBoost and KAMELEON-MLP models achieve high accuracy ( and respectively) and AUROC ( and respectively). Furthermore, for a positive class that, while less imbalanced, is still a minority, the Area Under the Precision-Recall Curve (AUPRC) is a vital metric. Here, the KAMELEON with sets the benchmark, significantly outperforming all other baselines. Unstructured models consistently yield the lowest performance for mortality prediction, AUROC values hover just above random chance (around ), and AUPRCs remain very low (max ), indicating a limited ability to discern between mortality and survival solely from clinical notes. SHAP results (Appendix Figure 8) demonstrate that lab results and vital signs strongly drive predictions, with prediction embeddings playing a less critical role compared to readmission. Overall, Table 2 shows that for both tasks, our multimodal model outperforms all individual structured and unstructured baselines across all metrics.
5.3 Ablation Study
To highlight the impact of reasoning features, we perform an ablation study (Table 8, Appendix). Removing reasoning features alone causes AUROC to drop from to for 30-day readmission. When we further remove demographics and prediction embeddings alongside reasoning, performance falls below , indicating that each component plays a complementary role. Figure 2 visualizes the contribution of reasoning, ’s prediction, and demographics to model performance. We also provide a similar plot for in-hospital mortality prediction in the Appendix. For this task, structured data played a more critical role, with less relative impact from reasoning and unstructured features compared to readmission prediction.
6 Conclusion
In this work, we introduced KAMELEON, a novel framework that effectively integrates multimodal EHR data, including structured clinical features and unstructured physician notes, enhanced by knowledge-augmented LLM reasoning, for robust clinical risk prediction. Our two-stage architecture demonstrated superior performance on both 30-day readmission and in-hospital mortality prediction, achieving a high AUROC score despite severely imbalanced datasets. KAMELEON consistently outperformed existing models, highlighting the synergy achieved by combining these diverse modalities.
This work underscores the significant potential of knowledge-augmented multimodal EHR modeling to enhance early intervention, optimize resource allocation, and improve patient care in complex clinical settings. Future work will focus on further validating KAMELEON’s generalizability across diverse clinical settings and exploring its application to a wider range of predictive healthcare tasks.
Scope and Promise for Social Impact. KAMELEON offers a strong opportunity to reduce avoidable hospital readmissions, a major driver of morbidity, cost, and financial penalties (Fry, Fluck, and Han 2021; Joynt and Jha 2013; Panagiotou et al. 2021; Zuckerman et al. 2017). The greatest impact is expected in high-risk groups, such as Outpatient Parenteral Antimicrobial Therapy (OPAT) patients, who face up to readmission rates (Wolie et al. 2024). OPAT programs are often underfunded and rely on manual, error-prone reviews. By identifying high-risk patients, our model enables focused use of limited resources, improving efficiency and outcomes.
We are exploring deployment in a large regional hospital for such high-impact populations, with later expansion to general readmissions.
@@ Appendix A Additional Details on Dataset
We use the Medical Information Mart for Intensive Care III (MIMIC-III) dataset (Johnson et al. 2016), a large-scale, publicly available critical care database comprising deidentified health-related data associated with over patients who were admitted to the intensive care units (ICUs) of Beth Israel Deaconess Medical Center between and . The dataset was curated to support clinical and epidemiological research and has since become a cornerstone resource in machine learning for healthcare.
@@ A.3 Untructured Data Representation for
Knowledge Graph Retrieval
@@
KG structure: This PubMed-derived knowledge graph supplements patient-specific data and improves reasoning by providing broader clinical context. The following triples represent extracted relations involving beta blocking agents from biomedical literature:
@@
Finding Similar Patients
@@ A.4 Full Example Prompt
Sample Inference Output
Table 7 shows a sample output generated by the fine-tuned LLM for the 30-day hospital readmission prediction task. Instead of returning a raw binary value (0 or 1), the model is prompted to generate both a prediction and its reasoning based on the patient’s diagnoses, procedures, medications, and comparisons with similar cases. This structured explanation allows the model to ground its prediction in clinical context, improving reliability and interpretability over naive classification. By incorporating rationale into the output, the LLM demonstrates better alignment with real-world clinical decision-making.
Patient ID: 25070_0
Patient ID: 93900_1
@@ Appendix B Additional results
Ablation on Readmission Table 8 presents an ablation study highlighting the performance impact of removing specific input features and components from models and for the readmission prediction task. The variations include removal of reasoning components, demonstrations, and predicted features.
KAMELEON achieves an accuracy of 0.80 and a Macro F1 score of 0.55, indicating a balanced performance across both classes. When the reasoning component is removed, we observe a spike in accuracy (0.92), but a sharp drop in recall for the positive class (Rec1 = 0.13), showing that the model becomes overly biased toward the negative class. This is further reflected in the reduced AUPRC and AUROC values.
Removing reasoning, demonstrations, and predicted features together leads to the worst performance (Macro F1 = 0.39), confirming that these components are critical for model generalization, especially in correctly identifying positive cases. The stepwise removal of features helps isolate the contribution of each component, underlining the importance of structured reasoning and prior knowledge in our approach.
Ablation on Mortality However, the SHAP values (Figure 8) show that the prediction embedding has a smaller impact compared to lab vitals, which contrasts with the readmission task. Additionally, we observe a higher AUROC for in-hospital mortality even without reasoning, exceeding . This high AUROC score suggests that predicting mortality is comparatively easier, likely due to the dataset being less imbalanced. Given access to lab vitals and diagnostic information, the model can more easily learn patterns associated with terminal illness.
| Model | Acc | Prec0 | Prec1 | Rec0 | Rec1 | Macro F1 | AUROC | AUPRC |
| our setup | 0.80 | 0.99 | 0.13 | 0.80 | 0.77 | 0.55 | 0.844 | 0.147 |
| without reasoning | 0.92 | 0.97 | 0.11 | 0.96 | 0.13 | 0.54 | 0.687 | 0.082 |
| without demo, reasoning, pred | 0.54 | 0.97 | 0.05 | 0.54 | 0.58 | 0.39 | 0.559 | 0.0433 |
| without reasoning, pred | 0.57 | 0.97 | 0.05 | 0.57 | 0.62 | 0.41 | 0.594 | 0.048 |
Analysis of distribution (In-Hospital Mortality)
Figure 9 shows predicted probabilities for in-hospital mortality within 30 days, separated by true class: survivors (blue) and deceased (red). The dashed line marks the optimal threshold (0.26) balancing sensitivity and specificity.
Class Separability Analysis Using t-SNE (In-hospital Mortality Task)
Figure 10 shows a two-dimensional t-SNE embedding of the test data, where the samples are colored by their true class labels (Class 0 in blue and Class 1 in red). The visualization reveals that the majority of the data forms a consistent, structured manifold dominated by Class 0 points, with only a small fraction of Class 1 points distributed across the embedding. Notably, a compact cluster of Class 1 samples appears on the right-hand side, indicating localized patterns that can be exploited by advanced models. This structure suggests that although Class 1 is relatively sparse, it exhibits distinct feature signatures in specific regions, which the proposed KAMELEON‑X model is designed to capture effectively, contributing to improved predictive performance.

### v2

@@ 4.1 Background
Problem statements. To demonstrate the effectiveness of our method, we study two popular clinical tasks: in-hospital mortality prediction and 30-day readmission prediction, e.g., [15, 1]. We define these problems formally after introducing some notation.
@@
30-day Readmission Problem. The objective is to determine whether the patient is readmitted to the hospital within 30 days following discharge from visit . We define the readmission indicator as:
@@
4.2 KAMELEON Framework
We propose a hybrid framework, KAMELEON, that integrates multimodal EHR data, including structured clinical components and unstructured physician notes (we will provide more details in Section 2.1), and external biomedical knowledge to predict two key clinical outcomes: in-hospital mortality and 30-day readmission. As shown in Figure 1, KAMELEON consists of two components: (1) an unstructured encoder (Section 4.2.1) that processes clinical notes and retrieves relevant biomedical knowledge using a PubMed-derived graph and knowledge-augmented reasoning, and outputs a prediction, along with its reasoning; and (2) a structured encoder (Section 4.2.2) that combines multiple time-series corresponding to vitals and tabular datasets (labs, medications, etc), along with the outputs from (i.e., both the prediction and the embedding associated with the reasoning it produces) with static features for downstream prediction. The notations used in Section 4.2, Algorithm 1, and 2 are explained in Table 4.
4.2.1 Unstructured Model,
@@
Generating Context. The first step of the framework is generating patient context. We use EHR data, including physicians’ notes, patient conditions, prescribed medications, and procedures. Since physician notes can be lengthy and exceed the context window of a small, locally fine-tuned LLM, we summarize them using an LLM. This approach addresses Challenge 2 (Long Context).
@@
Knowledge Graph Retrieval. To introduce domain-level medical knowledge, we build a biomedical knowledge graph (KG) by combining the Unified Medical Language System (UMLS)[29], PubMed abstracts, and large language model (LLM)-generated entity-relation triples. UMLS provides standardized biomedical concepts and relationships, and the entity-relation triples are structured facts extracted by the LLM in the form , capturing semantic connections. We apply the Leiden algorithm for community detection [45], which partitions the KG into semantically coherent and well-connected subgraphs. After clustering, we employ a separate LLM to generate a textual summary for each cluster. These summaries are produced by reasoning over the relationships among entities within each cluster, capturing the latent biomedical semantics encoded in the graph. Each summary serves as a high-level abstraction of the biomedical concepts and interactions within a subgraph. We embed these summaries using SentenceTransformer (MiniLM-L6-v2) [46] and retrieve the most relevant ones for each patient by computing semantic similarity with the embedded patient context. This process directly addresses Challenge 3 (Specialized Medical Domains) by enriching patient context with structured, domain-specific knowledge, improving the model’s understanding of specialized medical terminology.
@@
Finding Similar Patients. We provide additional context, by retrieving semantically similar patient visits using precomputed visit-level embeddings and a similarity index using the FAISS [47] library with inner-product search on L2-normalized embeddings, which effectively approximates cosine similarity. For each target patient visit, we retrieve the top 50 most similar patients while excluding self-matches and other visits from the same individual. Each retrieved patient is scored by similarity, and we filter them into positive and negative cohorts based on matching or non-matching ground truth labels (e.g., readmission vs. no readmission). The final output includes the top-k positive and negative similar patients (with k=1,2). Unlike KARE [15], we also provide the physician notes of the retrieved similar patients, enabling the language model to leverage more clinical context when assessing patient risk.
@@
Reasoning Module. In this module, we prepare inputs to fine-tune the LLM for clinical prediction. For each patient visit, we create a prompt with the patient’s context with the top-k most similar cases retrieved earlier. These similar cases guide the model by highlighting patterns in clinically comparable scenarios. We also add biomedical knowledge summaries from clustered subgraphs of a PubMed knowledge graph, providing literature-based context. Combining patient data, historical cases, and domain knowledge, we fine-tune the LLM to produce task-specific predictions with interpretable reasoning, supporting each outcome.
@@
Finetuning LLM. We fine-tune a LLaMA-3 8B model using the Unsloth framework [48, 49], which enables memory-efficient training via 4-bit quantization and Low-Rank Adaptation (LoRA) [50]. Prediction tasks are framed as instruction-following using Alpaca-style prompts with task description, patient context, and optional justification. Each prompt combines clinical notes, retrieved similar cases, and biomedical knowledge summaries. The model is trained via supervised learning to generate both predictions and reasoning. We use limited training steps with gradient accumulation and sequence lengths up to tokens. Unlike KARE, which trains larger models with higher compute, our method uses smaller, quantized models to reduce computational cost while maintaining interpretability and performance. Algorithm 1 outlines the training and inference procedures of . Additionally, an example prompt and its overall structure are provided in the Appendix Section A.3.
4.2.2 Structured Data Encoder,
@@
Time-varying variables. We extract hourly vitals and diagnoses in MIMIC-III, including heart rate, systolic and diastolic blood pressure, mean blood pressure, oxygen saturation, Glasgow Coma Scale (GCS) scores, glucose level, respiratory rate, temperature, weight, and pH.
Static metadata. Demographic and admission features include gender, ethnicity, admission type, location, insurance, language, religion, and hospital outcome.
Diagnoses, procedures, medications. ICD-9 codes, drug names are encoded via one-hot or counts. We compute binary indicators for key comorbidities (e.g., sepsis, infection, organ failure, dementia, cancer, diabetes).
Structured Data Preprocessing. The structured data is first transformed using a discretization step to enforce uniform temporal resolution and impute missing values. This is followed by normalization using precomputed mean and standard deviation statistics over the continuous variables.
Incorporating LLM Output. To augment the structured input, we include the LLM’s prediction and its tokenized reasoning. A Word2Vec [51] model is trained on this reasoning to generate embeddings. Each paragraph is represented by averaging its word vectors, producing a fixed-size vector. This reasoning vector is concatenated with the output label to form a feature vector for each visit.
@@ 4.2.3 Training
In our experiments, we benchmark several ML models for as KAMELEON-X, where X represents logistic regression, balanced random forests (BRF), LSTM, LightGBM, MLP, or XGBoost, selected for its effectiveness in capturing clinical patterns.
For MLP, we use weighted binary cross-entropy loss
@@
where and are the positive and negative class weights, respectively, used to address class imbalance by giving more emphasis to the minority class. We further use Synthetic Minority Over-sampling Technique (SMOTE) [52] to mitigate class imbalance, addressing Challenge 4 (Highly imbalanced data).
For KAMELEON-BRF model, which is designed to handle class imbalance. Here, each decision tree is trained on a bootstrapped sample drawn by under-sampling the majority class and combining it with all minority class examples, ensuring balanced class proportions at the tree level. Each split is chosen to minimize the Gini impurity:
where denotes teh proportion of class in a given node. The balanced sampling overcomes the issue with the dominance of the majority class and improves sensitivity to rare outcomes. Final predictions are obtained by aggregating probabilities across trees.
The complete training and inference pipeline for is outlined in Algorithm 2.
@@ Appendix A Additional Details on Dataset
We use the Medical Information Mart for Intensive Care III (MIMIC-III) dataset [35], a large-scale, publicly available critical care database comprising deidentified health-related data associated with over patients who were admitted to the intensive care units (ICUs) of Beth Israel Deaconess Medical Center between and . The dataset was curated to support clinical and epidemiological research and has since become a cornerstone resource in machine learning for healthcare.
@@ A.3.1 Generating Context
A.3.2 Knowledge Graph Retrieval
@@
KG structure: This PubMed-derived knowledge graph supplements patient-specific data and improves reasoning by providing broader clinical context. The following triples are an example how the relations are extracted, involving beta blocking agents from biomedical literature.
@@
A.3.3 Finding Similar Patients
@@ Appendix B Additional results
Patient ID: 9782_1
### Step 1: Analyzing Conditions 1. **Septicemia (except in labor)** - This is a severe condition that indicates systemic infection, which can lead to sepsis and high mortality risk. |
| 2. **Shock** - Indicates a critical state, often associated with poor outcomes. |
| 3. **Pneumonia** - This can be serious, especially in the context of septicemia and shock, as it can complicate respiratory function. |
| 4. **Schizophrenia and other psychotic disorders** - While this is a mental health condition, it does not directly impact mortality. |
| 5. **Deficiency and other anemia** - This can complicate other conditions, especially in the context of infection or respiratory issues. |
| 6. **Thyroid disorders** - Generally less critical unless there are acute complications. |
| ### Step 2: Analyzing Procedures 1. **Other vascular catheterization; not heart** - This is an invasive procedure that can carry risks, especially in the context of the patient’s other conditions. |
| 2. **No major surgeries or life-sustaining procedures** - The absence of life-sustaining interventions suggests that the patient may not be in immediate critical condition. |
| ### Step 3: Analyzing Medications 1. **Antibiotics (quinolone antibacterials, other antibacterials)** - Indicate the treatment of infection, which is critical given the septicemia and pneumonia. |
| 2. **Corticosteroids for systemic use** - Can be used in severe conditions but also carry risks, especially in the presence of infections. |
| 3. **Opioid analgesics** - Often used for pain management, which may indicate severe underlying conditions. |
| 4. **Cardiac stimulants** - Suggests potential heart issues. |
| 5. **High-risk medications** - The use of multiple medications indicates a complex health status, which increases the risk of adverse effects and complications. |
| ### Step 4: Comparing with Similar Patients 1. **Patient ID 15209_0** had conditions such as septicemia, respiratory failure, and shock, leading to a label of 0 (survival). This patient had similar severe conditions but also underwent extensive invasive procedures and had more acute conditions. |
| 2. **Patient ID 28061_0** had coronary atherosclerosis and cardiac issues leading to a label of 1 (mortality). This patient had significant cardiovascular issues and underwent multiple invasive procedures. |
| ### Step 5: Integrating Information |
| - The combination of septicemia, shock, and pneumonia in Patient ID 9782_1 is extremely concerning, especially considering the lack of major surgical interventions. The presence of multiple comorbidities (anemia, schizophrenia, and thyroid disorders) complicates the overall picture. |
| - The complexity and severity of the patient’s conditions (particularly septicemia and shock) indicate a high risk of mortality, but the absence of life-sustaining procedures suggests some resilience. |
| ### Conclusion |
| Given the severity and complexity of Patient ID 9782_1’s conditions, the presence of multiple high-risk medications, and the lack of immediate life-threatening interventions, the prediction leans towards survival (label 0). However, the critical nature of the patient’s conditions suggests that the risk of mortality is significant, and a careful monitoring plan is essential for managing the patient’s health. |
| ### Final Prediction |
| - **Prediction**: 0 (survival) |
Patient ID: 85258_1
### Step 1: Analyze Patient Conditions |
| 1. **Cancer of Esophagus**: This is a critical condition that is often associated with high mortality risk, particularly when combined with other severe conditions. |
| 2. **Secondary Malignancies**: The presence of secondary malignancies indicates that the patient has advanced cancer, further increasing the risk of mortality. |
| 3. **Pneumonia**: This is a severe infection that can lead to respiratory failure, especially in a patient with compromised respiratory function due to cancer. |
| 4. **Respiratory Failure**: The patient has been diagnosed with respiratory failure, indicating a critical state that significantly increases the risk of mortality. |
| 5. **Septicemia**: This is a serious systemic infection that can lead to multi-organ failure, which is a key indicator of mortality. |
| 6. **Deficiency Anemia**: This can complicate the patient’s overall health, particularly in the context of cancer and infection. |
| 7. **Other Comorbidities**: The patient has several other comorbidities, including disorders of teeth and jaw, white blood cell diseases, and fluid and electrolyte disorders, which compound the risk of mortality. |
| ### Step 2: Evaluate Procedures |
| 1. **Tracheostomy**: This is a significant surgical intervention that indicates severe respiratory compromise and is often performed in critically ill patients. |
| 2. **Mechanical Ventilation**: The patient is on mechanical ventilation, which is a life-sustaining measure typically reserved for critically ill patients. |
| 3. **Blood Transfusions**: The need for multiple blood transfusions suggests significant anemia or other severe conditions that require urgent intervention. |
| ### Step 3: Assess Medications |
| 1. **Broad Spectrum Antibiotics**: The patient is receiving broad-spectrum antibiotics, which are often used in the presence of sepsis or severe infections. |
| 2. **Corticosteroids**: The use of corticosteroids suggests a significant inflammatory response or severe underlying condition. |
| 3. **Antipsychotics and Antidepressants**: These medications suggest a complex mental health component, which can complicate treatment and recovery. |
| ### Step 4: Compare with Similar Patients |
| 1. **Patient ID: 28061_0**: This patient had a history of coronary artery disease and complications from medical care, leading to a mortality label of 1. The presence of critical cardiac issues and multiple comorbidities mirrors the complexity of Patient ID: 85258_1. |
| 2. **Patient ID: 14297_0**: This patient had several chronic conditions but ultimately had a label of 0, indicating that not all patients with complex conditions lead to mortality. The complexity of Patient ID: 85258_1, however, suggests a much higher risk. |
| ### Step 5: Synthesize Findings |
| - The combination of advanced cancer, severe infections (septicemia), respiratory failure, and invasive procedures (tracheostomy and mechanical ventilation) strongly indicates that the patient is in a critical state. |
| - The presence of multiple comorbidities and the complexity of the medications suggest that the patient is in a high-risk category for mortality. |
| ### Conclusion |
| Given the severity of the conditions, the invasive nature of procedures, and the complex medication regimen, the patient is at an extremely high risk of mortality. Therefore, the prediction for Patient ID: 85258_1 is: |
| **Prediction**: 1 (mortality) |
Appendix C Related Works
Clinical Prediction using Statistical and Machine Learning Methods: Structured data collected from electronic health records (EHRs) contain valuable information about patients and can be instrumental in capturing patient states and predicting medical outcomes (e.g., mortality) [15, 16, 17]. To leverage these data, traditional machine learning models—such as linear regression [8], decision trees [9], and Bayesian approaches [18]—use structured features as input to model the complex relationships between patient features and clinical outcomes. More recently, deep learning–based frameworks have also been explored to capture these complex dynamics and have shown great potential [10, 11, 12, 13, 14]. However, deep learning-based approaches remain largely restricted to structured data and rely heavily on statistical correlations between input features and outcome labels. Deep learning-based approaches also often lack interpretability grounded in explicit clinical knowledge and tend to be highly specific to the training dataset, limiting their generalizability to other clinical settings [15, 26].
NLP and Language Models: Unstructured clinical texts (e.g., physician notes, radiology reports, and nursing progress notes) contain rich information about patients [35, 53, 26]. Textual documentation includes clinicians’ reasoning and management plans that are not contained elsewhere in the EHR and which may provide direct and high-value insights. To harness this information, various natural language processing (NLP) methods have been proposed for predictive modeling [54, 55, 56]. Recently, large language models (LLMs) have demonstrated significant progress across a range of tasks, and researchers have begun exploring their potential for processing unstructured clinical text and supporting clinical reasoning [53, 28, 27]. However, one of the major challenges in applying LLMs to healthcare is the risk of errors and hallucinations [57, 22]. To address this, some works incorporate knowledge graphs to explicitly embed clinical knowledge and mitigate LLM inaccuracies [58, 26]. Despite these advances, integrating both structured EHR data and unstructured clinical text in a unified framework remains a challenging problem [59]. Existing approaches—such as prompt engineering [60], RAG [14], and data fusion [61, 62]—tend to be ad hoc and lack a systematic method for joint modeling.